OpenStudy (anonymous):

An airplane needs to reach a velocity of to take off. On a runway, what is the minimum acceleration necessary for the plane to take flight?

OpenStudy (mrnood):

we need the following info: minimum take off speed maximum take off roll (runway length) v^2 = u^2 +2as v is take off speed u = 0 a= acceleration s is take off distance (max take off roll or runway length)

OpenStudy (anonymous):

call runway length d then if starting from rest at one end with constant acceleration a we must reach speed s in distance d: d = (1/2) a t^2 s = a t or t = s/a, time rolling on runway so d = (1/2) a s^2/a^2 or d = (1/2) s^2/a so a = s^2/(2d)

OpenStudy (mrnood):

@damonc2 The variables oyu have used are not commonly used in this field. u or v0 is used for initial speed v or v1 used for final speed (s is often used for distance, as in my equations above) it is quite confusing to see s=at. the full equation is v = u +at (but u=0 in this case)